# Gaussian approximation on the Skorokhod space via Malliavin calculus and regularization

### v1

@@ 1. Introduction
where the matrix $(\Gamma(F,-L^{-1}F)=(\Gamma(F_{i},-L^{-1}F_{j}))_{1\leq i\leq j\leq n}$ is given by $\Gamma(F_{i},-L^{-1}F_{j}))=\left\langle DF_{i},-DL^{-1}Fj\right\rangle_{\mathfrak{H}}$, with $D$ the Malliavin derivative, $L^{-1}$ the pseudo-inverse of the generator and $\mathfrak{H}$ the Hilbert space where the underlying isonormal Gaussian process is defined. The pioneering and pathbreaking work [NP09] of Nourdin and Peccati in dimension one was generalized to multiple but finite dimension by the same authors and Reveillac shortly afterwards (see [NPR10]). For a book-length treatment of the finite-dimensional theory, we refer to  [NP12], its far reaching ramifications in diverse areas of mathematics are well-known and classical.
@@ 2.7. Wiener-Itô chaos and multiple integrals in Banach spaces
called the $E$-valued multiple Wiener-Itô integral of order $p$, whose range is $\mathcal{H}_{p}(E)$. The isometry property of $I_{p}^{\mathbb{R}}$ is only retained in the form of an equivalence of norms: For any $s\in[1,\infty)$ and any $p\in\mathbb{N}_{0}$ there exist positive constants $\mathfrak{c}_{p,s}(E)$ and $\mathfrak{C}_{p,s}(E)$ such that for all $T\in\gamma^{p}_{sym}(\mathfrak{H},E)$ one has
(14) $\mathfrak{c}_{p,s}(E)\left\lVert T\right\rVert_{\gamma^{p}(\mathfrak{H},E)}\leq\left\lVert I_{p}^{E}(T)\right\rVert_{L^{s}(\Omega;E)}\leq\mathfrak{C}_{p,s}(E)\left\lVert T\right\rVert_{\gamma^{p}(\mathfrak{H},E)},$
@@
(15) $\mathfrak{c}_{p,s}=2^{\frac{p-1}{s}}\sqrt{\frac{p!}{p^{p}}}\kappa_{s,2}^{-1}(E)\qquad\text{and}\qquad\mathfrak{C}_{p,s}=4^{\frac{3p+1}{s}}\sqrt{\frac{p^{p}}{p!}}\kappa_{s,2}(E),$
where $\kappa_{s,t}(E)$ denotes the Kahane-Khintchine constant for the $L^{s}-L^{t}$-norm comparison of Gaussian (or Rademacher) sums (see for example [HvVW17, Thm. 6.2.6] for a precise definition). In the special case where $E$ is a Hilbert space and $s=2$, (14) of course reduces to the classical Wiener-Itô isometry, i.e. the inequalities (14) become equalities, which hold for $\mathfrak{c}_{p,2}(E)=\mathfrak{C}_{p,2}(E)=\sqrt{p!}$.
@@ 3. The carré du champ operator and integration by parts formulæ
$\qquad=\mathbb{E}\left(\mathbb{E}^{\prime\prime}\left(\left\lVert\sum_{\mathbf{u}\in[n]^{r},\mathbf{s}\in[n]^{p-r}}W^{\prime\prime}_{\mathbf{u}}W_{\mathbf{s}}\,x_{(\mathbf{u},\mathbf{s})}\right\rVert_{\pi}\,\left\lVert\sum_{\mathbf{v}\in[n]^{r},\mathbf{t}\in[n]^{q-r}}W^{\prime\prime}_{\mathbf{v}}W^{\prime}_{\mathbf{t}}\,y_{(\mathbf{u},\mathbf{t})}\right\rVert_{\pi}\right)^{2}\right),$
@@
Let $E$ and $F$ be Banach spaces, $X\in\mathbb{D}^{1,2}(E)$ and $Y\in\mathbb{D}^{1,2}(F)$. Then
@@
We first treat the case where $X\in\mathcal{S}(\mathbb{R})\otimes E$ and $Y\in\mathcal{S}(\mathbb{R})\otimes F$ of the form
@@
$=\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},y_{k}\right)_{F^{\ast},F}\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)$
$=\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)\otimes y_{k}\right)_{F^{\ast},F},$
@@
$\operatorname{tr}_{\pi}\left(Q_{f(X),Y}\right)$ $=\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},\sum_{k=1}^{n}\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)\otimes y_{k}\right)_{F^{\ast},F}$
$=\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},-DL^{-1}Y\mathfrak{h}_{l}\right)_{F^{\ast},F}$
$=\sum_{l}b_{f^{\prime}}\left(DX\,\mathfrak{h}_{l},-DL^{-1}Y\mathfrak{h}_{l}\right)$
$=\left(b_{f^{\prime}},\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}.$
@@
For the general case, we approximate $X\in\mathbb{D}^{1,2}(E)$ and $Y\in\mathbb{D}^{1,2}(F)$ by sequences $(X_{n})_{n\in\mathbb{N}}\subseteq\mathcal{S}(\mathbb{R})\otimes E$ and $(Y_{n})_{n\in\mathbb{N}}\subseteq\left(\mathcal{S}(\mathbb{R})\otimes F\right)$, such that $\left\lVert X-X_{n}\right\rVert_{1,2}\to 0$ and $\left\lVert Y-Y_{n}\right\rVert_{1,2}\to 0$. Then,
$\operatorname{tr}_{\pi}\left(Q_{f(X_{n}),Y_{n}}\right)=\left(b_{f^{\prime}(X)},\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}.$
@@
$\mathbb{E}$ $\left(\left|\left(Y-\mathbb{E}(Y),f(X)\right)_{F,F^{\ast}}-\left(Y_{n}-\mathbb{E}(Y_{n}),f(X_{n})\right)_{F,F^{\ast}}\right|\right)$
$\leq\mathbb{E}\left|\left(Y-Y_{n},f(X)\right)_{F,F^{\ast}}\right|+\mathbb{E}\left|\left(\mathbb{E}(Y-Y_{n}),f(X)\right)_{F,F^{\ast}}\right|+\mathbb{E}\left|\left(Y_{n}-\mathbb{E}(Y_{n}),f(X)-f(X_{n})\right)_{F,F^{\ast}}\right|$
$\leq\left\lVert f\right\rVert_{\infty}\left(2\left\lVert Y-Y_{n}\right\rVert_{L^{1}(\Omega;F)}+\left\lVert Y_{n}-\mathbb{E}(Y_{n})\right\rVert_{L^{2}(\Omega;F)}\left\lVert X-X_{n}\right\rVert_{L^{2}(\Omega;E)}\right),$
which tends to zero as $n\to\infty$. Furthermore
@@
$\qquad\leq\mathbb{E}\left|\left(f^{\prime}(X),\Gamma_{\pi}(X,-L^{-1}Y)-\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right)_{\pi}\right|+\mathbb{E}\left|\left(f^{\prime}(X)-f^{\prime}(X_{n}),\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right)_{\pi}\right|$
$\qquad\leq\left\lVert f^{\prime}\right\rVert_{\infty}\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right\rVert_{\pi}\right)$
(22) $\qquad\qquad\qquad+\mathbb{E}\left(\left\lVert f^{\prime}(X)-f^{\prime}(X_{n})\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}\,\left\lVert\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right\rVert_{\pi}\right).$
The first expectation in (22) vanishes in the limit due to Lemma 3.1 and Lemma 2.5, as
$\mathbb{E}$ $\left(\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right\rVert_{\pi}\right)$
$\leq\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X-X_{n},-L^{-1}Y)\right\rVert_{\pi}\right)+\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X_{n},-L^{-1}(Y-Y_{n}))\right\rVert_{\pi}\right)$
$\leq\left\lVert D(X-X_{n})\right\rVert_{L^{2}(\Omega;\gamma(\mathfrak{H},E)}\,\left\lVert DL^{-1}Y\right\rVert_{L^{2}(\Omega;\gamma(\mathfrak{H},F))}$
$\qquad\qquad+\left\lVert DX_{n}\right\rVert_{L^{2}(\Omega;\gamma(\mathfrak{H},E))}\,\left\lVert DL^{-1}(Y_{n}-Y)\right\rVert_{L^{2}(\Omega;\gamma(\mathfrak{H},F))}$
$\leq\left(\left\lVert DX\right\rVert_{L^{2}(\Omega;\gamma(\mathfrak{H},E))}+\left\lVert DX_{n}\right\rVert_{L^{2}(\Omega;\gamma(\mathfrak{H},E))}\right)\left\lVert D(Y_{n}-Y)\right\rVert_{L^{2}(\Omega;\gamma(\mathfrak{H},F))}.$
For the second summand in (22), note that the above calculation and the assumption that $DX_{n}\to DX$ and $DY_{n}\to DY$ in probability implies $L^{1}(\Omega)$- and almost sure convergence of $\left\lVert\Gamma(X_{n},-L^{-1}Y_{n})\right\rVert_{\pi}$. Furthermore, the sequence of random variables $\left\lVert f^{\prime}(X)-f^{\prime}(X_{n})\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}$ is bounded (by $2\left\lVert f^{\prime}\right\rVert_{\infty}$) and converges to zero in probability (by the open mapping theorem, using $X_{n}\to X$ in probability and $f^{\prime}$ continuous). Therefore, the sequence $(G_{n})$ of random variables defined by
$G_{n}=\left\lVert\Gamma(X_{n},-L^{-1}Y_{n})\right\rVert_{\pi}\,\left\lVert f^{\prime}(X)-f^{\prime}(X_{n})\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}$
is uniformly integrable and converges in probability to zero. The Vitali Convergence Theorem hence implies that $G_{n}\to 0$ in $L^{1}(\Omega)$.
@@ 5. Contraction bounds on Wiener chaos
We now extend the product formula (see for example [Nua06, Prop. 1.1.3] or  [NP12, Thm. 2.7.10]) to the vector-valued setting. As in the previous sections, given $u\in E$ and $v\in F$, we denote by $u\otimes v$ both the simple tensor in $E\otimes F$ and the rank one operator in $\mathcal{F}(F1G^{\ast},E)$. Which definition is meant can always be inferred from the context.
@@
(31) $I_{p}(f)\otimes I_{q}(f)=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r,\pi}g\right)$
@@
$I_{p}(f)\otimes I_{q}(f)=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r}g\right)$
@@
$\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-Q_{X,Y}\right\rVert_{L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)}\\ \leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}(E\widehat{\otimes}_{\pi}F)\,\left\lVert f\widetilde{\otimes}_{r,\pi}g\right\rVert_{\gamma^{p+q-2r}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)},$
@@
$\left\lVert\Gamma(X,-L^{-1}Y)-R_{X,Y}\right\rVert_{L^{2}(\Omega;\mathfrak{N}(F^{\ast},E))}\\ \leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}(\mathfrak{N}(F^{\ast},E))\,\left\lVert f\widetilde{\otimes}_{r}g\right\rVert_{\gamma^{p+q-2r}(\mathfrak{H},\mathfrak{N}(F^{\ast},E))},$
where the positive constants $\mathfrak{C}_{m,2}(\cdot)$ and $a_{p,q,r}$ are defined at (15) and (32), respectively.
@@
$\qquad\qquad\leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}(E\widehat{\otimes}_{\pi}F)\,\left\lVert f\widetilde{\otimes}_{r,\pi}g\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$
$\qquad\qquad\leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}(E\widehat{\otimes}_{\pi}F)\,\left\lVert f\otimes_{r,\pi}g\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$

### v2

@@ 1. Introduction
where the matrix $\Gamma(F,-L^{-1}F)=(\Gamma(F_{i},-L^{-1}F_{j}))_{1\leq i\leq j\leq n}$ is given by $\Gamma(F_{i},-L^{-1}F_{j})=\left\langle DF_{i},-DL^{-1}Fj\right\rangle_{\mathfrak{H}}$, with $D$ the Malliavin derivative, $L^{-1}$ the pseudo-inverse of the generator and $\mathfrak{H}$ the Hilbert space where the underlying isonormal Gaussian process is defined. The pioneering and pathbreaking work [NP09] of Nourdin and Peccati in dimension one was generalized to multiple but finite dimension by the same authors and Reveillac shortly afterwards (see [NPR10]). For a book-length treatment of the finite-dimensional theory, we refer to  [NP12], its far reaching ramifications in diverse areas of mathematics are well-known and classical.
@@ 2.7. Wiener-Itô chaos and multiple integrals in Banach spaces
called the $E$-valued multiple Wiener-Itô integral of order $p$, whose range is $\mathcal{H}_{p}(E)$. The isometry property of $I_{p}^{\mathbb{R}}$ is only retained in the form of an equivalence of norms: For any $s\in[1,\infty)$ and any $p\in\mathbb{N}_{0}$ there exist positive constants $\mathfrak{c}_{p,s}$ and $\mathfrak{C}_{p,s}$ such that for all $T\in\gamma^{p}_{sym}(\mathfrak{H},E)$ one has
(14) $\mathfrak{c}_{p,s}\left\lVert T\right\rVert_{\gamma^{p}(\mathfrak{H},E)}\leq\left\lVert I_{p}^{E}(T)\right\rVert_{L^{s}(\Omega;E)}\leq\mathfrak{C}_{p,s}\left\lVert T\right\rVert_{\gamma^{p}(\mathfrak{H},E)},$
@@
(15) $\mathfrak{c}_{p,s}=2^{\frac{p-1}{s}}\sqrt{\frac{p!}{p^{p}}}\kappa_{s,2}^{-1}\qquad\text{and}\qquad\mathfrak{C}_{p,s}=4^{\frac{3p+1}{s}}\sqrt{\frac{p^{p}}{p!}}\kappa_{s,2},$
where $\kappa_{s,t}$ denotes the Kahane-Khintchine constant for the $L^{s}-L^{t}$-norm comparison of Gaussian (or Rademacher) sums (see for example [HvNVW17, Thm. 6.2.6] for a precise definition). In the special case where $E$ is a Hilbert space and $s=2$, (14) of course reduces to the classical Wiener-Itô isometry, i.e. the inequalities (14) become equalities, which hold for $\mathfrak{c}_{p,2}=\mathfrak{C}_{p,2}=\sqrt{p!}$.
@@ 3. The carré du champ operator and integration by parts formulæ
$\qquad=\mathbb{E}\left(\mathbb{E}^{\prime\prime}\left(\left\lVert\sum_{\mathbf{u}\in[n]^{r},\mathbf{s}\in[n]^{p-r}}W^{\prime\prime}_{\mathbf{u}}W_{\mathbf{s}}\,x_{(\mathbf{u},\mathbf{s})}\right\rVert_{E}\,\left\lVert\sum_{\mathbf{v}\in[n]^{r},\mathbf{t}\in[n]^{q-r}}W^{\prime\prime}_{\mathbf{v}}W^{\prime}_{\mathbf{t}}\,y_{(\mathbf{u},\mathbf{t})}\right\rVert_{F}\right)^{2}\right),$
@@
Let $E$ and $F$ be Banach spaces, $X\in\mathbb{D}^{1,2}(E)$ and $Y\in\mathbb{D}^{1,2}(F)$. Then,
@@
We first treat the case where $X$ and $Y$ are of the form
@@
$=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},y_{k}\right)_{F^{\ast},F}\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)\right)$
$=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)y_{k}\right)_{F^{\ast},F}\right),$
@@
$\operatorname{tr}_{\pi}\left(Q_{f(X),Y}\right)$ $=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},\sum_{k=1}^{n}\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)y_{k}\right)_{F^{\ast},F}\right)$
$=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},-DL^{-1}Y\mathfrak{h}_{l}\right)_{F^{\ast},F}\right)$
$=\mathbb{E}\left(\sum_{l}b_{f^{\prime}(X)}\left(DX\,\mathfrak{h}_{l},-DL^{-1}Y\mathfrak{h}_{l}\right)\right)$
$=\mathbb{E}\left(\left(b_{f^{\prime}(X)},\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}\right).$
@@
$\operatorname{tr}\left(R_{f(X),Y}\right)=\operatorname{tr}\left(\mathbb{E}\left(f^{\prime}(X)\Gamma(X,-L^{-1}Y)\right)\right)=\mathbb{E}\left(\operatorname{tr}\left(f^{\prime}(X)\Gamma(X,-L^{-1}Y)\right)\right),$
where the second equality follows from the fact that the Bochner integral commutes with bounded linear mappings.
For the general case, we approximate $X\in\mathbb{D}^{1,2}(E)$ and $Y\in\mathbb{D}^{1,2}(F)$ by sequences $(X_{n})_{n\in\mathbb{N}}\subseteq\mathcal{F}\left(\mathcal{S}(\mathbb{R}),E\right)$ and $(Y_{n})_{n\in\mathbb{N}}\subseteq\mathcal{F}\left(\mathcal{S}(\mathbb{R}),F\right)$, such that $\left\lVert X-X_{n}\right\rVert_{1,2}\to 0$ and $\left\lVert Y-Y_{n}\right\rVert_{1,2}\to 0$. Possibly passing to subsequences, we assume that $(X_{n})_{n\in\mathbb{N}}$ and $(Y_{n})_{n\in\mathbb{N}}$ converge almost surely. Then,
$\operatorname{tr}_{\pi}\left(Q_{f(X_{n}),Y_{n}}\right)=\mathbb{E}\left(\left(b_{f^{\prime}(X_{n})},\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right)_{\pi}\right).$
@@
$\operatorname{tr}_{\pi}$ $\left(Q_{f(X),Y}\right)-\operatorname{tr}_{\pi}\left(Q_{f(X_{n}),Y_{n}}\right)$
$=\mathbb{E}\left(\left|\left(Y-\mathbb{E}(Y),f(X)\right)_{F,F^{\ast}}-\left(Y_{n}-\mathbb{E}(Y_{n}),f(X_{n})\right)_{F,F^{\ast}}\right|\right)$
$\leq\mathbb{E}\left|\left(Y-\mathbb{E}(Y),f(X)-f(X_{n})\right)_{F,F^{\ast}}\right|+\mathbb{E}\left|\left(Y-Y_{n},f(X_{n})\right)_{F,F^{\ast}}\right|+\mathbb{E}\left|\left(\mathbb{E}(Y-Y_{n}),f(X_{n})\right)_{F,F^{\ast}}\right|$
$\leq 2\left\lVert f\right\rVert_{\infty}\mathbb{E}\left(\left\lVert Y-Y_{n}\right\rVert_{F}\right)+2\mathbb{E}\left(\left\lVert Y\right\rVert_{F}\right)\mathbb{E}\left(\left\lVert f(X)-f(X_{n})\right\rVert_{F^{\ast}}\right)$
which tends to zero as $n\to\infty$ by dominated convergence and the assumption that $Y_{n}\to Y$ almost surely. Furthermore
@@
$\qquad\leq\mathbb{E}\left|\left(f^{\prime}(X)-f^{\prime}(X_{n}),\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}\right|$
$\qquad\qquad\qquad+\mathbb{E}\left|\left(f^{\prime}(X_{n}),\Gamma_{\pi}(X,-L^{-1}Y)-\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right)_{\pi}\right|$
$\qquad\leq\mathbb{E}\left(\left\lVert f^{\prime}(X)-f^{\prime}(X_{n})\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}\,\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)\right\rVert_{\pi}\right)$
$\qquad\qquad\qquad+\left\lVert f^{\prime}\right\rVert_{\infty}\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right\rVert_{\pi}\right)$
and both terms on the right hand side vanish in the limit due to dominated convergence, as $\left\lVert\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})-\Gamma_{\pi}(X,-L^{-1}Y)\right\rVert_{\pi}$ converges to zero by the boundedness of $\Gamma_{\pi}$ and $\left\lVert f^{\prime}(X)-f^{\prime}(X_{n})\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}$ converges to zero by continuity of $f$ and the almost sure convergence of $(X_{n})_{n\in\mathbb{N}}$.
@@ 5. Contraction bounds on Wiener chaos
We now extend the product formula (see for example [Nua06, Prop. 1.1.3] or  [NP12, Thm. 2.7.10]) to the vector-valued setting. As in the previous sections, given $u\in E$ and $v\in F$, we denote by $u\otimes v$ both the simple tensor in $E\otimes F$ and the rank one operator in $\mathcal{F}(G^{\ast},E)$. Which definition is meant can always be inferred from the context.
@@
(30) $I_{p}(f)\otimes I_{q}(g)=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r,\pi}g\right)$
@@
$I_{p}(f)\otimes I_{q}(g)=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r}g\right)$
@@
$\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-Q_{X,Y}\right\rVert_{L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)}\\ \leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\widetilde{\otimes}_{r,\pi}g\right\rVert_{\gamma^{p+q-2r}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)},$
@@
$\left\lVert\Gamma(X,-L^{-1}Y)-R_{X,Y}\right\rVert_{L^{2}(\Omega;\mathfrak{N}(F^{\ast},E))}\\ \leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\widetilde{\otimes}_{r}g\right\rVert_{\gamma^{p+q-2r}(\mathfrak{H},\mathfrak{N}(F^{\ast},E))},$
where the positive constants $\mathfrak{C}_{m,2}$ and $a_{p,q,r}$ are defined at (14) and (31), respectively.
@@
$\qquad\qquad\leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\widetilde{\otimes}_{r,\pi}g\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$
$\qquad\qquad\leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\otimes_{r,\pi}g\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$